OpenStudy (anonymous):

. Tommy gives baseball cards to four of his friends. He gives Kiana 3 cards less than he gives Gary. He gives Monty 7 cards more than he gives Kiana. He gives Willa twice as many as he gives Monty. How many cards does Tommy give to each friend if he gives out a total of 74 cards?

OpenStudy (anonymous):

Let's call these friends by their first initial: K = Kiana, G = Gary, M = Monty, W = Willa Now we can convert the question to a system of linear equations: Gary has 3 cards more than Kiana, so G = K + 3 Monty has 7 cards more than Kiana, so M = K + 7 Willa has twice as many cards as Monty, so W = 2M = 2(K + 7) = 2K + 14 74 cards given out = all the friends combined, so 74 = K + G + M + W = K + (K + 3) + (K + 7) + (2K + 14) = 5K + 3 + 7 + 14 = 5K + 24 50 = 5K 10 = K, so Kiana had 10 cards G = K + 3 = 13, so Gary had 13 cards M = K + 7 = 17, so Monty had 17 cards W = 2M = 34, so Willa had 34 cards Checking our work, 10 + 13 + 17 + 34 = 74 cards total.